OpenStudy (anonymous):

Using the given points and line, determine the slope of the line. (-3, 0) and (2, 7) PLEASE HELP!

OpenStudy (anonymous):

Slope formula: \[m = \frac{ y2 - y1 }{ x2 - x1 }\] All we have to do is plug the points in and solve. (-3, 0) and (2, 7) y2 = 7 y1 = 0 x2 = 2 x1 = -3 \[m = \frac{ 7 - 0 }{ 2 - (-3)}\] 7- 0 = 7 2 - (-3) = 2 + 3 = 5 \[m = \frac{ 7 }{ 5 }\]

OpenStudy (anonymous):

Thank you!! @BunnyBree

OpenStudy (anonymous):

You're welcome! :D